the length of a rectangle is 3 inches more than twice its width. its area is 65 inches sq. find the length and width of the rectangle

Answer by mananth(16949)

let width =x
length = 2x+3
x(2x+3)= 65
2x^2+3x=65
2x^2+3x-65=0
2x^2+13x-10x-65=0
2x(x-6.5)-10(x-6.5)=0
(2x-10)(x-6.5)=0
x= 5 & 6.5
The length and breadth 5 & 13